The length of a rectangular piece of land is 60 yards more than two times its width. The perimeter is 540 yards. Find its dimensions.
Answer by addingup(3677)

The perimeter:
2L+2W = 540
And the problem says:
L = 2W+60 So let's substitute in the equation above:
2(2W+60)+2W = 540
4W+120+2W = 540
6W = 420
W = 70 And the length:
70(2)+60 = 200
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
2L+2W = 540
2(200)+2(70) = 540
400+140 = 540 We have the correct answer